# Efficient Training-Free Online Routing for High-Volume Multi-LLM Serving

### v1

@@ B.1 Discussion on Our Theoretical Assumptions
Random arrival order. We assume that the queries arrive in a random order, i.e., they can be picked adversarially, but their order is randomly permuted. Note that this is a weaker requirement than assuming the queries are sampled i.i.d. from an unknown distribution. This is because if the queries are sampled i.i.d., then they also satisfy the random order requirement, and thus our assumption is weaker. Furthermore, the random order model is also a popular assumption in many other online algorithms [39, 11], since it typically allows one to go beyond worst-case hardness and obtain meaningful theoretical guarantees. In practice, we perform robustness studies by changing the query order and observe that our algorithm still outperforms baselines; see Section˜5.1.
@@
Our assumption is further corroborated by the relaxation used in Section˜3.1, where if $\frac{\hat{C}_{opt}}{s_{max}}$ is sufficiently large, the original MILP can be closely approximated by its fractional LP relaxation with only a negligible optimality gap. We validate this in our main experiments over three benchmarks with the observed optimality gaps being: (i) 0.016% on SPROUT, (ii) 0.086% on RouterBench, and (iii) 0.3% on Open LLM Leaderboard v2.
@@ B.3 Proof of Lemma 3
$\sum_{i}z_{i}\leq O\left(\epsilon^{2}\sqrt{(1+\delta)C_{opt}{\hat{C}}/\alpha}\right)$
@@ B.5 Proof of Lemma 5
From Equation˜5, it follows immediately that $\sum_{i}F_{i}(\gamma^{*})=F(\gamma^{*})$.
@@ B.6 Proof of Theorem 1
Appendix C More Experiment Results
@@ C.4 Total Budget
C.5 Historical Data Size & ANNS/KNN Candidates
@@
C.6 Detailed Ablation Study
@@
C.7 Routing Analysis by Query Types
@@
Appendix D Extended Related Work
@@
The second paradigm leverages historical or auxiliary datasets to estimate the performance and cost of each query and select the optimal LLM accordingly [43, 50, 23, 13, 37, 27, 48, 52, 15, 20, 56]. In this line of work, several approaches train model-based predictors using these labeled historical datasets. For example, RouterLLM [43] trains a BERT-base or causal LLM on the historical data annotated with human preferences to enhance routing accuracy. CARROT [50] trains Roberta-base models to estimate both performance and cost for each query, enabling selection of the optimal LLM under any user-defined trade-off between quality and cost. HybridLLM [13] adopts the synthetic preference labels from the MixInstruct dataset [26] based on BARTScore [61], and trains a single BERT-based router for query routing. Zooter [37] distills reward signals from training queries and applies tag-based label enhancement to train a routing function that facilitates the expert selection. TensorOpera [52] introduces a soft label-based strategy based on the BERT similarity scores to train the BERT-based predictor for query routing. GraphRouter [15] constructs a heterogeneous graph to capture the contextual relationship between the query requirements and the LLM capabilities for informed routing decisions. While these training-based approaches are effective for certain LLM deployment settings, they introduce nontrivial training overhead. Furthermore, adapting them to varying LLM deployment configurations typically requires retraining, which makes them unsuitable for resource-constrained online routing with diverse LLM deployment configurations.

### v2

@@ B.1 Discussion on Our Theoretical Assumptions
Random arrival order. We assume that the queries arrive in a random order, i.e., they can be picked adversarially, but their order is randomly permuted. Note that this is a weaker requirement than assuming the queries are sampled i.i.d. from an unknown distribution. This is because if the queries are sampled i.i.d., then they also satisfy the random order requirement, and thus our assumption is weaker. Furthermore, the random order model is also a popular assumption in many other online algorithms [39, 11], since it typically allows one to go beyond worst-case hardness and obtain meaningful theoretical guarantees. In practice, we perform robustness studies by changing the query order and observe that our algorithm still outperforms baselines; see Section˜4.1.
@@
Our assumption is further corroborated by the relaxation used in Section˜2.1, where if $\frac{\hat{C}_{opt}}{s_{max}}$ is sufficiently large, the original MILP can be closely approximated by its fractional LP relaxation with only a negligible optimality gap. We validate this in our main experiments over three benchmarks with the observed optimality gaps being: (i) 0.016% on SPROUT, (ii) 0.086% on RouterBench, and (iii) 0.3% on Open LLM Leaderboard v2.
@@ B.3 Proof of Lemma 3
$\sum_{i}z_{i}\leq O\left(\epsilon^{2}\sqrt{(1+\delta)C_{opt}{\hat{C}}/\alpha}\right).$
@@ B.5 Proof of Lemma 5
From Section˜2.2, it follows immediately that $\sum_{i}F_{i}(\gamma^{*})=F(\gamma^{*})$.
@@ B.6 Proof of Theorem 1
Appendix C Additional Experimental Results
@@ C.4 Total Budget
Table 7: Per-query routing latency (ms) of different algorithms on RouterBench under varying numbers of queries.
Algorithm 4000 queries 6000 queries 8000 queries 10000 queries
Greedy-Perf 0.203 0.203 0.201 0.203
Greedy-Cost 0.233 0.230 0.226 0.233
KNN-Perf 0.478 0.475 0.469 0.475
KNN-Cost 0.300 0.282 0.281 0.278
BatchSplit 0.300 0.282 0.281 0.278
Ours 0.060 0.060 0.060 0.064
Table 8: Routing performance on RouterBench under two challenging historical-data conditions: noisy data label and out-of-distribution (OOD) data. Perf represents the Performance, PPC represents Performance per Cost, and Tput represents Throughput.
Noisy Data Label Out-of-Distribution (OOD)
Algorithm Perf Cost PPC Tput Perf Cost PPC Tput
Random 1403.250 0.419 3350.430 3227 1495.35 0.518 2888.14 3308.00
Greedy-Perf 1323.600 0.370 3573.750 2471 807.15 0.252 3204.76 1643.00
Greedy-Cost 1660.200 0.461 3604.720 4047 1677.45 0.561 2989.30 4044.00
KNN-Perf 1323.950 0.369 3585.390 2472 816.10 0.254 3212.31 1658.00
KNN-Cost 1660.200 0.461 3604.720 4047 1679.70 0.561 2993.23 4047.00
BatchSplit 1757.250 0.456 3856.680 3974 1828.55 0.526 3474.42 4283.00
Roberta-Perf 83.000 0.045 1836.330 84 9.50 0.017 559.15 10.00
Roberta-Cost 358.000 0.075 4787.593 1616 1692.65 0.562 3013.92 4037.00
Ours 2319.650 0.440 5221.280 4775 2050.8 0.532 3851.53 4644.00
C.5 Routing Overhead & Quality of Historical Data
Routing Overhead. We evaluate the routing latency of Algorithm˜1 under varying query volumes. We compare Algorithm˜1 with training-free routing baselines, focusing on the core online routing stage that follows the embedding step. This focus is motivated by two key reasons: (1) The embedding step can operate asynchronously and is fully decoupled from the core routing logic, where its outputs are shareable across retrieval, logging, and other downstream tasks. (2) The embedding model is interchangeable (e.g., MiniLM, GTE-small, bge-small), while our core contribution lies in the core routing algorithm itself. As shown in Table 7, our method incurs negligible decision overhead compared to existing training-free baselines. Even under high query volume (10,000 queries), it introduces only 0.064ms of routing latency per query and consistently achieves the lowest decision overhead (except the random routing strategy, which exhibits significantly worse performance).
Quality of Historical Data. To assess robustness to data quality, we consider a noisy setting with label perturbations and an out-of-distribution (OOD) setting with distribution shift. In the noisy setting, we introduce two types of strong noise simultaneously to the data label. For each performance label $d_{ij}$, we randomly flip its value with a probability of 20%. For each cost label $g_{ij}$, we apply multiplicative, mean-preserving noise consisting of: (i) log-normal jitter $\left(\log\mathcal{N}(0,\sigma^{2})-\frac{1}{2}\sigma^{2}\right)$ with $\sigma=0.25$, and (ii) a spike factor of 3.0 applied with 2% probability. In the OOD setting, we construct a distribution shift using RouterBench, which contains 13 diverse datasets. Specifically, we use MMLU as the historical dataset and evaluate on the remaining 12 datasets, which differ significantly in data distribution. Results in Table 8 show that our algorithm consistently outperforms all baselines under both noisy label and OOD settings. It achieves the highest overall performance, the best performance-per-cost ratio, and the greatest throughput in both settings, demonstrating strong robustness to label noise and distribution shift.
C.6 Historical Data Size & ANNS/KNN Candidates
@@
C.7 Detailed Ablation Study
@@
C.8 Routing Analysis by Query Types
@@
Appendix D Query-Traffic Considerations in LLM Routing
In the practical LLM-serving systems, an additional important routing factor is query traffic patterns. However, the primary goal of this work is to develop a cost-efficient online routing algorithm, a problem that has gained increasing attention in recent literature [43, 23, 50], especially given the high computational cost of LLMs. To isolate this core algorithmic problem, we adopt a standard assumption widely used in the online algorithms literature [11, 39, 19]: incoming queries arrive in a randomly permuted order. This assumption, discussed in detail in Appendix B.1, allows us to abstract away complex traffic dynamics and focus on the fundamental cost-efficiency routing problem.
Under this assumption, our algorithm is designed to operate on query-intrinsic features that are independent of time and query order, such as response performance score and cost per query across the models. In contrast, operational-state features, such as current queue lengths or system load, are beyond the scope of our current formulation and are more aligned with a related but orthogonal problem: load balancing under traffic constraints.
Still, we believe our algorithm offers a natural foundation for incorporating such traffic-aware considerations. One potential approach is to discretize time into short windows and track the per-model backlog of unserved workload. Within each time window, a corresponding "time-bucket" MILP can be solved with additional traffic constraints, and any unserved workload can be carried over to the next window. For instance, let ${\ell}_{ij}$ denote the service time of query $j$ on LLM $i$, $b_{it}$ be the backlog of unserved workload for LLM $i$ at the start of time window $t$, and $A_{it}$ denote the available service capacity for LLM $i$ during window $t$. Then, for each window $t$ and LLM $i$, we can add the following constraint to the MILP:
$\sum_{j}{\ell}_{ij}x_{ij}\leq\max\{0,A_{it}-b_{it}\},\quad\forall i,t$ (8)
This “time-bucket” formulation helps smooth bursts and reduce queuing delays. We leave the integration of such traffic-aware routing mechanisms to future work.
Appendix E Extended Related Work
@@
The second paradigm leverages historical or auxiliary datasets to estimate the performance and cost of each query and select the optimal LLM accordingly [43, 50, 23, 13, 37, 27, 48, 52, 15, 20, 56]. In this line of work, several approaches train model-based predictors using these labeled historical datasets. For example, RouterLLM [43] trains a BERT-base or causal LLM on the historical data annotated with human preferences to enhance routing accuracy. CARROT [50] trains Roberta-base models to estimate both performance and cost for each query, enabling selection of the optimal LLM under any user-defined trade-off between quality and cost. HybridLLM [13] adopts the synthetic preference labels from the MixInstruct dataset [26] based on BARTScore [61], and trains a single BERT-based router for query routing. Zooter [37] distills reward signals from training queries and applies tag-based label enhancement to train a routing function that facilitates expert selection. TensorOpera [52] introduces a soft label-based strategy based on the BERT similarity scores to train the BERT-based predictor for query routing. GraphRouter [15] constructs a heterogeneous graph to capture the contextual relationship between the query requirements and the LLM capabilities for informed routing decisions. While these training-based approaches are effective for certain LLM deployment settings, they introduce nontrivial training overhead. Furthermore, adapting them to varying LLM deployment configurations typically requires retraining, which makes them unsuitable for resource-constrained online routing with diverse LLM deployment configurations.
@@
Appendix F Broader Impacts
Our algorithm proposes an efficient, training-free solution for online routing in high-volume, budget-constrained LLM serving. It significantly reduces serving costs and improves performance, which enables more sustainable, scalable, and cost-effective multi-LLM deployment. As a technical contribution, it does not pose direct societal risks. We believe it promotes broader accessibility to LLM services, and future work can further address potential concerns by incorporating responsible use guidelines.